How is a wave group formed ? Obtain an expression for the velocity of a wave group. What is the relation connecting the group and phase velocities ?
1
Expert's answer
2019-05-17T11:49:53-0400

The velocity with which the overall envelope shape of the wave's amplitudes is called as the velocity of a wave group. The individual wavelets of differing wavelengths traveling at different speeds form a wave group.


We consider a wave packet as a function of position x and time t: α(x,t).

Let A(k) be its Fourier transform at time t = 0,

"α(x,0)= \int_{ -∞}^ {∞} dk A(k)e^{ ikx } (1)"

where ω is implicitly a function of k.


By the superposition principle, the wavepacket at any time t is



"α(x,t)= \int_{ -∞}^ {∞} dk A(k)e ^{ i(kx-ωt) } (2)"

where ω is implicitly a function of k.

After integrating (2) we got:


"v_g = \frac {δω }{δk} (3)"



where ω is the wave's angular frequency (usually expressed in radians per second), and k is the angular wavenumber.


Phase velocity vp and group velocity vg are related through Rayleigh's formula,


"v_g=v_p(1-\frac {ω }{v_p} \frac {dv_p }{dω})^{-1} (4)"